the perimeter of an equalateral triangle is 15 inches. find the length of the altitude.
Answer by checkley77(12569)

An equilateral triangle has equal sides
15/3=5 in. for each side.
Now we have a right triangle with a base=5/2 & a hypotenuse=5.
The third side is the altitude (b).
a^2+b^2=c^2
(5/2)^2+b^2=5^2
25/4+b^2=25
b^2=25-25/4
b^2=25-6.25
b^2=18.75
b=sqrt18.75
b=4.33 in. ans. for the altitude.